OpenStudy (aaaa):

Write the solution for the following problem as an ordered pair: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work. medals given

OpenStudy (anonymous):

y+2x=46 y=x+4

OpenStudy (anonymous):

x+4+2x=46 3x+4=46 3x=42 x=14 y=x+4 y=14+4=18